# Pre-trained Universal Medical Image Transformer

### 12 Dec 2023 v1

@@ Abstract
Self-supervised learning has emerged as a viable method to leverage the abundance of unlabeled medical imaging data, addressing the challenge of labeled data scarcity in medical image analysis. In particular, masked image modeling (MIM) with visual token reconstruction has shown promising results in the general computer vision (CV) domain and serves as a candidate for medical image analysis. However, the presence of heterogeneous 2D and 3D medical images often limits the volume and diversity of training data that can be effectively used for a single model structure. In this work, we propose a spatially adaptive convolution (SAC) module, which adaptively adjusts convolution parameters based on the voxel spacing of the input images. Employing this SAC module, we build a universal visual tokenizer and a universal Vision Transformer (ViT) capable of effectively processing a wide range of medical images with various imaging modalities and spatial properties. Moreover, in order to enhance the robustness of the visual tokenizer’s reconstruction objective for MIM, we suggest to generalize the discrete token output of the visual tokenizer to a probabilistic soft token. We show that the generalized soft token representation can be effectively integrated with the prior distribution regularization through a constructive interpretation. As a result, we pre-train a universal visual tokenizer followed by a universal ViT via visual token reconstruction on 55 public medical image datasets, comprising over 9 million 2D slices (including over 48,000 3D images). This represents the largest, most comprehensive, and diverse dataset for pre-training 3D medical image models to our knowledge. Experimental results on downstream medical image classification and segmentation tasks demonstrate the superior performance of our model and improved label efficiency. Our code is available at https://github.com/function2-llx/PUMIT.
Introduction
Over the last decade, medical image analysis, a field crucial for disease diagnosis and treatment planning, has notably benefited from the development of deep learning and shown success in various tasks (Litjens et al. 2017). Training these deep learning models requires a substantial amount of labeled data, while the availability of labeled medical data is often limited due to the expensive acquisition process. To address this issue, there has been a growing interest in self-supervised learning approaches, which learn representations from the large amounts of unlabeled data by designing self-supervised pretext tasks (Shurrab and Duwairi 2022). Despite the abundance of unlabeled data, the medical images are highly heterogeneous in terms of imaging modalities (CT, MRI, ultrasound, etc.) and, especially, spatial properties (2D and 3D spatial dimension, various voxel spacing and spatial shape), making it non-trivial to process all kinds of medical images with a unified model structure. The mainstream models are commonly designed to handle images of a single spatial property for a single model structure (Isensee et al. 2021). As a result, most of the previous works only leverage medical images data of similar spatial properties, limiting the amount and diversity of data for pre-training.
Recent advancements of the Vision Transformer (ViT) (Dosovitskiy et al. 2021) in the field of computer vision offer a promising solution to this challenge. The ViT converts the 2D or 3D spatial structure of an image into 1D patch sequences using a patch embedding layer and processes these sequences with transformer blocks. It has shown great potential in being used as a universal model to handle a wide range of medical images with diverse spatial properties. A recent work, UniMiSS (Xie et al. 2022a), has made initial attempts in this direction, leveraging both 2D and 3D unlabeled medical images for self-supervised pre-training. The authors designed a switchable patch embedding (SPE) module composed of both 2D and 3D patch embedding layers, which adaptively switches to either layer depending on the input dimension. While UniMiSS represents a significant step forward in breaking the “dimensionality barrier”, its patch embedding layer does not adaptively handle 3D medical images of various voxel spacings.
Following the success in the natural language processing (NLP) domain, masked image modeling (MIM) with ViT has emerged as a powerful self-supervised approach for learning vision representations. During MIM pre-training, an image is divided into non-overlapping patches (Dosovitskiy et al. 2021), with a subset being randomly masked out. The ViT model is tasked with predicting information related to the masked patches based on the visible ones. Currently, MIM pretext tasks are primarily categorized into three approaches: 1) pixel (voxel) value regression (Xie et al. 2022b), 2) visual token reconstruction with a pre-trained visual tokenizer (Bao et al. 2022), and 3) vision feature prediction (Wei et al. 2022). While pixel value regression is straightforward for MIM, it tends to focus on low-level information, given the inherent sparsity of image information (He et al. 2022; Fang et al. 2023b). Although the other two approaches of MIM can effectively capture high-level semantics, their application to medical image analysis is limited by the lack of universal models.
In this work, we aim to bridge the aforementioned gap by pre-training a universal visual tokenizer followed by a universal ViT with MIM via visual token reconstruction on a large volume of medical images. To achieve this, we introduce a spatially adaptive convolution (SAC) module, which adaptively adjusts convolution parameters based on the voxel spacing of the input image. Models built with the SAC module are capable of effectively processing a wide range of medical images with various imaging modalities and spatial properties. We also propose several improvements to the visual tokenizer to enhance the robustness of the reconstruction objective for MIM. A visual tokenizer is typically implemented using a vector-quantized variational autoencoder (VQ-VAE) (van den Oord, Vinyals, and Kavukcuoglu 2017). However, it can suffer from issues of severe codebook collapse and low codebook utilization with deterministic quantization, or the perturbations to the reconstruction process introduced by stochastic quantization during pre-training (Zhang et al. 2023). To address these issues, we generalize the discrete token to the soft token, which is represented as a categorical distribution over the codebook. We demonstrate the prior distribution regularization (Zhang et al. 2023) can be generalized to such soft token representation through a constructive interpretation. Furthermore, we extend this technique by maximizing the KL divergence of individual token distributions from the prior distribution, refining the sharpness of the learned token distributions. The contributions of our work are summarized as follows:
  • We propose the spatially adaptive convolution module that serves as a building block of the universal models for medical images. Consequently, we build such universal visual tokenizer and ViT with it for pre-training.
  • We generalize the discrete token in the visual tokenizer to the probabilistic soft token, mitigating codebook collapse without stochastic quantization. The token distributions are learned to be both diverse and sharpened through the extended prior distribution regularization.
@@ Introduction
  • We train the visual tokenizer followed by the ViT on 55 public medical image datasets, with over 9 million 2D slices (including over 48,000 3D images). This represents the largest, most comprehensive, and diverse dataset for pre-training 3D medical image models to our knowledge.
@@
  • Our pre-trained models are fine-tuned on downstream medical image classification and segmentation tasks. Experimental results demonstrate the superior performance of our model and improved label efficiency.
Related Works
Self-supervised Learning on Medical Images
Self-supervised learning (SSL) aims to learn meaningful representations from unlabeled data by formulated pretext tasks. The application of SSL in medical image analysis can be mainly classified into the following three categories (Shurrab and Duwairi 2022). Contrastive SSL is a widely used technique that commonly models image similarity and dissimilarity through a family of data augmentations (Azizi et al. 2021; Sowrirajan et al. 2010; He et al. 2023). Predictive SSL aims to learn robust representations by formulating the pretext task as a classification or regression problem, such as geometric transformation prediction (Zhou et al. 2023), rotation angle prediction (Zhu et al. 2020; Tang et al. 2022), and context reconstruction (Zhou et al. 2021a; Chen et al. 2019). Generative SSL enables models to learn representations by either reconstructing input data or generating new examples from input distribution, commonly with autoencoders (Zhou et al. 2021b) and GANs (Liu et al. 2021). Recent works also explore leveraging auto-generated volumetric shapes for 3D medical image segmentation (Tadokoro, Yamada, and Kataoka 2023).
@@
Inspired by the masked language modeling (MLM) in NLP (Devlin et al. 2019), masked image modeling (MIM) with transformer-based models has emerged as a viable SSL approach in the CV domain. Some representative works include SimMIM (Xie et al. 2022b) and MAE (He et al. 2022) with pixel value regression, BEiT family (Bao et al. 2022; Peng et al. 2022) and iBOT (Zhou et al. 2022) with visual token reconstruction, and MVP (Wei et al. 2022) and EVA series (Fang et al. 2023b, a) with vision feature prediction. Recent works also explore learning multimodal representations with MIM (Baevski et al. 2022, 2023) and investigate the scaling behaviors of MIM (Xie et al. 2023). However, there have been fewer attempts to apply MIM to medical image analysis (Chen et al. 2023), largely due to the significant differences between medical images and natural images.
Universal Models for Medical Images
A universal model capable of processing medical images with diverse properties under a unified model structure is crucial for leveraging a large amount of unlabeled data available for SSL. One straightforward approach to build such a universal model is to use a 2D model to process multiple 2D slices split from a 3D volume (Nguyen et al. 2023b, a). However, using a 2D model also requires additional mechanism to address the cross-plane contextual information inherent in 3D images. UniMiSS (Xie et al. 2022a) presents an alternative approach based on the ViT model, which converts the input 2D or 3D image into sequences of 1D image patches using a switchable patch embedding (SPE) layer. These sequences are then processed by transformer blocks. Additionally, the weight inflation technique (Zhang et al. 2022), which adapts pre-trained weights from 2D models to 3D models, also provides meaningful insights into the development of universal models for medical images.
Preliminary
In this section, we briefly review the concept of the visual tokenizer. Similar to a tokenizer in NLP, a visual tokenizer can convert an image $x$ into a $D\times H\times W$ grid of visual tokens, where each token belongs to a finite set $\mathcal{V}$, referred to as the visual codebook. Such a visual tokenizer can be implemented using a VQ-VAE (van den Oord, Vinyals, and Kavukcuoglu 2017), a variant of VAE (Kingma and Welling 2013) where the latent variables follow discrete categorical distributions over $\mathcal{V}$. Specifically, an encoder learns a posterior distribution $q(z\mid x)$ to output a $D\times H\times W$ grid of categorical token distributions over $\mathcal{V}$ given the image $x$. Token samples can be drawn from these categorical distributions in a differentiable manner using Gumbel-Softmax reparameterization (Maddison, Mnih, and Teh 2017; Jang, Gu, and Poole 2017) and straight-through estimators (Bengio, Léonard, and Courville 2013). Each token in $\mathcal{V}$ corresponds to a $d$-dimensional embedding vector from an embedding space $\bm{e}\in\mathbb{R}^{|\mathcal{V}|\times d}$. A decoder learns to reconstruct the original image given the embeddings of the drawn token samples. The prior distribution of the latent variable at each position is the uniform distribution over $\mathcal{V}$. Similarly to vanilla VAE, such visual tokenizer can be trained by maximizing the evidence lower bound (ELBO).
Methodology
Spatially Adaptive Convolution
Voxel spacing is an important spatial property of 3D medical images, representing the physical size of a voxel. Conventional image-processing networks, such as convolutional networks, operate on the voxel grid but disregard the voxel spacing information. To address this discrepancy, a common strategy is to resample the images in a dataset to a uniform spacing (Isensee et al. 2021). However, this approach is impractical for pre-training on a large volume of images with diverse voxel spacings.
Instead, we propose the adaptive adjustment of the convolution parameters based on the voxel spacing of the input. Typically, a 3D medical image consists of a stack of multiple slices, with voxel spacing defined by two degrees of freedom: the slice thickness $s_{\mathrm{slice}}$ and the isotropic in-plane spacing $s_{\mathrm{plane}}$. A 2D image can be viewed as a special case with only one slice and sufficiently large slice thickness. To explain how the SAC module works, we define the degree of anisotropy (DA) of a feature map to be $\mathrm{DA}=\max\left\{0,\left\lfloor\log_{2}\frac{{s_{\mathrm{slice}}}}{s_{\mathrm{plane}}}\right\rfloor\right\}$. The SAC module is essentially a 3D convolution or transposed convolution initialized with kernel sizes and strides that are isotropic across all dimensions, with the following three variants:
  • Downsampling: Both convolution kernel size and stride are $2^{k}$, where $k$ is a non-negative integer. The convolution weight along the depth dimension is reduced to the size of $2^{\max\{k-\mathrm{DA},0\}}$ through sum pooling, and the stride of the depth dimension is similarly adjusted to $2^{\max\{k-\mathrm{DA},0\}}$. The spacing of the output feature map is multiplied by the adjusted stride.
  • $3^{3}$ convolution: The convolution kernel size is 3 and stride is 1 on the depth dimension. If $\mathrm{DA}>0$, the convolution weight is reduced along the whole depth dimension to the size of 1 by sum pooling.
  • Upsampling: Similar to downsampling, both transposed convolution kernel size and stride are $2^{k}$, where $k$ is a non-negative integer, but the reduced kernel size and stride of the depth dimension is $2^{\min\{k,\mathrm{DA}_{0}-\mathrm{DA}\}}$, where $\mathrm{DA}_{0}$ is the DA of the input image. The spacing of the output feature map is divided by the adjusted stride.

### 24 Jan 2024 v2

# Building Universal Foundation Models for Medical Image Analysis with Spatially Adaptive Networks

@@ Abstract
Recent advancements in foundation models, typically trained with self-supervised learning on large-scale and diverse datasets, have shown great potential in medical image analysis. However, due to the significant spatial heterogeneity of medical imaging data, current models must tailor specific structures for different datasets, making it challenging to leverage the abundant unlabeled data. In this work, we propose a universal foundation model for medical image analysis that processes images with heterogeneous spatial properties using a unified structure. To accomplish this, we propose spatially adaptive networks (SPAD-Nets), a family of networks that dynamically adjust the structures to adapt to the spatial properties of input images, to build such a universal foundation model. We pre-train a spatial adaptive visual tokenizer (SPAD-VT) and then a spatial adaptive Vision Transformer (SPAD-ViT) via masked image modeling (MIM) on 55 public medical image datasets. The pre-training data comprises over 9 million image slices, representing the largest, most comprehensive, and most diverse dataset to our knowledge for pre-training universal foundation models for medical image analysis. The experimental results on downstream medical image classification and segmentation tasks demonstrate the superior performance and label efficiency of our model. Our code is available at https://github.com/function2-llx/PUMIT.
1 Introduction
Figure 1: The first two stages of U-Net encoders designed by nnU-Net for Task 1 (left) and Task 5 (right) of MSD challenge Antonelli et al. (2022), and the version of our proposed SPAD-Nets (middle). Convolution parameters ($k$ for kernel size, $s$ for stride) along the depth dimension are indicated. The structures designed by nnU-Net for datasets with different spacing have incompatible parts (marked as red). The SPAD-Nets are able to handle images from both tasks by adapting structures to input spatial properties.
Conventional deep learning models require a substantial amount of labeled data to be trained, but the availability of labeled data is often limited in domains such as health and medicine due to the expensive acquisition process. Moreover, most of these learning models are developed in a task-specific manner, where individual models are designed and trained on a narrow range of data, further limiting the applications of models. Recently, the field of artificial intelligence (AI) has witnessed a paradigm shift from training task-specific models to developing foundation models, which are typically trained on large-scale and diverse datasets using self-supervised learning (SSL), and can be applied to various downstream tasks via fine-tuning or prompting Bommasani et al. (2021). In general vision domain, for instance, without the need for expensive annotations, foundation models can be pre-trained on raw images with masked image modeling (MIM) and achieve promising results in downstream vision tasks such as classification, detection, and segmentation He et al. (2022); Fang et al. (2023b).
However, general visual foundation models cannot be applied to medical images directly due to the significant differences between natural images and medical images. While most natural images are 2D and of RGB modality, medical images are diverse in terms of imaging modalities such as X-ray, ultrasound, CT, and MRI, and dimensions as both 2D and 3D images are commonly used. Although 2D and 3D convolutional neural networks (CNNs) have been used for 2D and 3D images, respectively, their model structures are inherently incompatible with each other due to different spatial dimensions. Previous works attempted to break such “dimensionality barrier” by switching to Vision Transformers (ViTs) Dosovitskiy et al. (2021); Xie et al. (2022), but CNNs still play indispensable roles in ViTs as their patch embedding layers. On the other hand, 3D medical images that are made up of stacked 2D slices can have a wide range of voxel spacing, that is, the physical size of a voxel, commonly represented by an in-plane spacing $s_{\mathrm{plane}}$ and a slice thickness $s_{\mathrm{slice}}$. When the ratio between $s_{\mathrm{slice}}$ and $s_{\mathrm{plane}}$ ($\nicefrac{{s_{\mathrm{slice}}}}{{s_{\mathrm{plane}}}}$) is overly large, the image content may change substantially across slices. Neural networks typically process images based on voxel grids, which are assumed to have isotropic voxel spacing, without considering the actual physical size of voxels. As a result, most mainstream models adapt their structures to align with the voxel spacing statistics of individual datasets Isensee et al. (2021). Yet, such practice can make model structures incompatible among datasets with different voxel spacing statistics, as illustrated in Figure 1. As a result, developing foundation models for medical image analysis is challenging because of the lack of unified model structures capable of handling medical imaging data with heterogeneous spatial properties.
In this work, we develop a universal foundation model for a broad range of medical images, including both 2D and 3D images with multiple imaging modalities and diverse voxel spacing. To accomplish this, we introduce the concept of spatially adaptive networks (SPAD-Nets), a family of networks that dynamically adjust their structures to adapt to the spatial properties of input images. SPAD-Nets are implemented by replacing the building blocks of conventional networks, such as convolutions, with spatially adaptive ones. Specifically, we propose spatially adaptive convolution (SPAD-conv) to serve as the building blocks of SPAD-Nets. A SPAD-conv shares the same set of parameters as a conventional 3D convolution, and its parameters are dynamically transformed to adapt to the input. For instance, given an input image with a large $\nicefrac{{s_{\mathrm{slice}}}}{{s_{\mathrm{plane}}}}$ value, a SPAD-conv will reduce its convolution kernel size and stride along the depth dimension accordingly. With models built with SPAD-Nets, we pre-train a spatially adaptive visual tokenizer (SPAD-VT) and then a spatially adaptive ViT (SPAD-ViT) via masked image modeling (MIM) on large-scale public medical image datasets. The contributions of our work are summarized as follows:
  • We propose SPAD-Nets, a family of networks used for building universal foundation models for medical image analysis. Based on SPAD-Nets, we manage to develop the SPAD-ViT, which is, to our knowledge, the first universal foundation model for medical image analysis.
@@ 1 Introduction
  • We train a spatially adaptive visual tokenizer (SPAD-VT) followed by a SPAD-ViT on 55 public medical image datasets with over 9 million image slices. This is the largest, most comprehensive, and most diverse dataset to our knowledge for pre-training universal foundation models for medical image analysis.
@@
  • Our pre-trained models are fine-tuned on downstream medical image classification and segmentation tasks. The experimental results demonstrate the superior performance and label efficiency of our model.
2 Related Works
Universal Models for Medical Image Analysis
A universal model capable of processing medical images with diverse spatial properties using a unified model structure is crucial for leveraging a large amount of unlabeled data available when developing foundation models. One straightforward approach to building such a universal model is to use a 2D model to process multiple 2D slices split from a 3D volume Nguyen et al. (2023). However, using a 2D model also requires additional mechanism to address the cross-plane contextual information inherent in 3D images. UniMiSS Xie et al. (2022) presents an alternative approach based on the ViT model, which converts the input 2D or 3D image into a sequence of 1D image patches using a switchable patch embedding (SPE) layer. The sequence is then processed by transformer blocks. Additionally, the weight inflation technique Zhang et al. (2022) is also proposed to adapt pre-trained weights from 2D models to 3D models, which inspires the development of SPAD-Nets.
@@ 2 Related Works
Inspired by the masked language modeling (MLM) in NLP Devlin et al. (2019), masked image modeling (MIM) with transformer-based models has emerged as a viable SSL approach in the CV domain. Some representative works include MAE He et al. (2022) with pixel value regression, BEiT family Bao et al. (2022) with visual token reconstruction, and EVA series Fang et al. (2023b, a) with vision feature prediction. However, there have been fewer attempts to apply MIM to medical image analysis Chen et al. (2023), largely due to the significant differences between medical images and natural images.
3 Preliminary
In this section, we briefly review key concepts about spatial properties in medical images. We then discuss how these concepts influence the design of current mainstream models for individual tasks of medical image analysis, and how they complicate the development of a foundation model.
A 3D medical image, such as a CT scan, is a stack of cross-sectional 2D image slices of a human body generated by a medical imaging device Buzug (2011). Similar to a typical digital image that can be represented by a 2D array of pixel values, each slice contains a 2D intensity array, and arrays of all slices are stacked to form the 3D array of voxel intensities for the whole 3D image. The spatial dimensions within the slice plane are referred to as the in-plane dimensions, and the dimension along which slices are stacked is referred to as the depth dimension. A 3D medical image contains information on voxel spacing, which is the physical size of a voxel along each spatial dimension and is represented as $s_{\mathrm{slice}}\times s_{\mathrm{h}}\times s_{\mathrm{w}}~{}(\mathrm{mm}^{3})$. Here, $s_{\mathrm{slice}}$ is the slice thickness (or equivalently, the distance between adjacent slices), $s_{\mathrm{h}}$ and $s_{\mathrm{w}}$ are spacings along in-plane dimensions. In our work, we treat 2D images as a special case of 3D images with only one slice and a sufficiently large $s_{\mathrm{slice}}$. Moreover, all images in this work have an isotropic in-plane spacing, $s_{\mathrm{h}}=s_{\mathrm{w}}$, which we denote as $s_{\mathrm{plane}}$.
Figure 2: Illustration of our pre-training framework for both SPAD-VT and SPAD-ViT. The models built with our proposed SPAD-Nets can process a wide range of images using a unified model structure. Note that SPAD-VT is trained first (optimizing $L_{\mathrm{VT}}$), and is fixed during the training of SPAD-ViT (optimizing $L_{\mathrm{MIM}}$).
The voxel spacing is crucial for accurate and comprehensive feature extraction from a medical image, such as measuring the physical volume of an object or the distance between objects. However, general neural networks only operate on the voxel grids of the intensity array that are assumed to have isotropic voxel spacing but disregard the actual physical size of voxels. To adapt a model to a dataset, nnU-Net Isensee et al. (2021), the widely used approach for medical image segmentation, proposes the following two empirical strategies: (1) Resampling all images to the same target spacing, such as the median value of spacings of all training images. (2) Adapting the network structure when the voxel spacing is significantly anisotropic, that is, the ratio between $s_{\mathrm{slice}}$ and $s_{\mathrm{plane}}$ is far from (mostly, much larger than) $1$. Take the downsampling stage in the U-Net encoder illustrated in Figure 1 as an example: the feature map will not be downsampled along the depth dimension at a stage if the value of $\nicefrac{{s_{\mathrm{slice}}}}{{s_{\mathrm{plane}}}}$ is larger than $2$.
These strategies have been validated to be effective for individual tasks by nnU-Net with extensive experiments and largely adopted by later works in 3D medical image analysis Zhou et al. (2023); Brugnara et al. (2023). However, both strategies become infeasible when developing a foundation model on large-scale and diverse datasets due to the following reasons: 1) Resampling all images from diverse datasets to the same target spacing will result in significant information loss and resampling artifacts, which can destroy delicate details in medical images. 2) The network structures adapted for individual tasks are static and can be incompatible with other datasets, whereas a unified structure is essential to leverage the large-scale data together. In conclusion, it is critical to develop new strategies to address the heterogeneous spatial properties of medical images that are applicable across diverse datasets for the development of foundation models for medical image analysis.
4 Methodology
4.1 Overview
Our solution to the challenge described in Section 3 is to design a universal model for a broad range of medical images with diverse spatial properties, built with our proposed SPAD-Nets. At each stage of the model, the structure can be dynamically adjusted to the one that is adapted to the input. We are then able to devise a pre-training framework based on masked image modeling (MIM) with visual token reconstruction Bao et al. (2022), illustrated in Figure 2. For an input image, a pre-trained visual tokenizer converts it to a sequence of tokens, where each token represents a patch on the image. With random patches on the image masked out, the training objective is to predict tokens of masked patches given the visible patches. As there is no pre-trained visual tokenizer for medical images available, especially a universal one, we also focus on the development of a visual tokenizer first.
4.2 SPAD-Nets
SPAD-Nets are implemented by replacing the building blocks of conventional networks with spatially adaptive ones. Specifically, we propose the spatially adaptive convolution (SPAD-conv) to serve as the convolutional building blocks of SPAD-Nets. In general, a SPAD-conv shares the same set of parameters as a conventional 3D convolution, which is termed as its base network and has isotropic convolution kernel sizes and strides across all spatial dimensions. When encountering an input with anisotropic spacing, the parameters are transformed accordingly to adapt to the input before applying to it. The design idea of SPAD-conv is largely inspired by and modified from the empirical rules of designing a static network structure adapting to images of a dataset proposed by nnU-Net. Given an input feature map, let $s_{\mathrm{plane}}$ and $s_{\mathrm{slice}}$ be the in-plane spacing and slice thickness as introduced in Section 3, respectively. The general principle for designing SPAD-conv can be summarized as follows: (1) if $s_{\mathrm{slice}}<2s_{\mathrm{plane}}$, the anisotropy of the input is considered to be tolerable and the input can be directly processed by the base network; (2) otherwise, when $s_{\mathrm{slice}}\geq 2s_{\mathrm{plane}}$, the input is considered to be overly anisotropic where changes across slices can be significant, and the base network must adjust the structure to reduce the information aggregation across the depth dimension. Based on this principle, we design three types of SPAD-conv: downsampling, k3s1, and upsampling, whose base networks are used as fundamental components in common vision models, such as U-Net. We now detail these types of SPAD-conv with an example of turning a conventional U-Net into a SPAD-Nets counterpart.
Downsampling
In the U-Net encoder, the feature map is typically downsampled with a ratio of 2 at each stage through a convolution with a stride of 2 and a kernel size of 2 or 3. When using the downsampling type of SPAD-conv with the same convolution as base network, given an overly anisotropic input ($s_{\mathrm{slice}}\geq 2s_{\mathrm{plane}}$), the downsampling along the depth dimension will be disabled. Specifically, the weight of convolution kernel along the depth dimension is reduced to the size of $1$ through sum pooling, and the stride along the depth dimension is similarly adjusted to $1$. The voxel spacing of the output feature map is multiplied by the adjusted stride.
K3s1
The convolution with a kernel size of $3$ and a stride of 1, abbreviated as k3s1 by us, is used in both the U-Net encoder and decoder for feature extraction. When using the k3s1 type of SPAD-conv with the same convolution as base network, given an overly anisotropic input ($s_{\mathrm{slice}}\geq 2s_{\mathrm{plane}}$), the feature aggregation across the depth dimension is disabled by reducing the weight of convolution kernel along the depth dimension to the size of 1 through sum pooling.
Upsampling
In the U-Net decoder, the feature map is typically upsampled with a ratio of 2 at each stage through a transposed convolution with a stride of 2 and a kernel size of 2 or 3. When using the upsampling type of SPAD-conv with the same transposed convolution as base network, however, the feature map is upsampled along the depth dimension if and only if it is downsampled along the depth dimension within the corresponding symmetric stage in the encoder. This is because the upsampling is considered to be a reverse process of downsampling in terms of spatial properties.